# Discovering Large-Scale Structure at in the C3VO Survey

### v1

@@ Abstract
The Charting Cluster Construction with VUDS and ORELSE (C3VO) survey is an ongoing imaging and spectroscopic campaign aiming to map out the growth of structure up to and was born from the combination of the VIMOS Ultra Deep Survey (VUDS) and the Observations of Redshift Evolution in Large-Scale Environments (ORELSE) survey. As we previously accomplished with the ORELSE survey, we apply our technique known as Voronoi tessellation Monte-Carlo (VMC) mapping to search for serendipitous galaxy overdensities at in the three C3VO fields. We also apply the same technique to mock observations of simulated galaxies with properties derived from the GAlaxy Evolution and Assembly (GAEA) semi-analytic model (SAM) in order to judge the effectiveness of our as a function of redshift, total mass, and fraction of spectroscopic redshifts. We find typical completeness and purity values of the order 30-50%, with a strong dependence on mass and redshift, with values as high as 80% and 70%, respectively, in the best-case scenario for . In the C3VO fields, we were able to recover many of the previously known structures in the literature as well as find hundreds of new overdensity candidates, once again demonstrating the powerful capabilities of VMC mapping when applied to wide-field optical and infrared galaxy evolution surveys at ever higher redshifts.
@@ 1 Introduction
Galaxy groups and clusters are among the largest gravitationally bound structures in the Universe. These large-scale structures at different redshifts provide a means of studying cosmology as well as the effects of environment on galaxy evolution at various stages in time. Protoclusters are the progenitors of galaxy clusters, which at some point will gravitationally collapse and become virialized. Clusters may be distinguished from field galaxies through a number of methods, such as the identification of a well-defined red sequence or a hot intracluster medium (ICM). Red sequence techniques to find clusters were first employed by the Red Sequence Cluster Survey (RCS; Gladders & Yee, 2000, 2005). The RCS made use of two filters, one in the optical and one in the near-infrared, to identify galaxy overdensities by features such as Balmer absorption or the 4000Å break, which would be present in the filter wavelengths out to redshifts of . With the addition of mid-infrared filters, red sequence searches can be extended up (e.g., Andreon et al., 2009; Muzzin et al., 2009; Wilson et al., 2009; Galametz et al., 2012; Cooke et al., 2015; Noirot et al., 2016; Golden-Marx et al., 2019).
@@
Various cluster surveys at intermediate redshift, such as the Observations of Redshift Evolution in Large-Scale Environments (ORELSE; Lubin et al., 2009) survey, Gemini Cluster Astrophysics Spectroscopic Survey (GCLASS; Muzzin et al., 2012), and the Gemini Observations of Galaxies in Rich Early ENvironments (GOGREEN; van der Burg et al., 2020) survey, have demonstrated how the scaling relations between color, density, morphology, and the star formation rate (SFR) change up to relative to the local universe. Studies from these surveys have shown that quenching is likely driven by internal mass-based mechanisms as well as processes that are at play in overdense environments. There are claims that these processes are largely separable for galaxy samples at (e.g., Peng et al., 2010; Kovač et al., 2014; Guglielmo et al., 2015), but some studies argue otherwise (e.g., De Lucia et al., 2012). The scaling relations weaken or reverse as redshift increases, and the two modes of quenching may also no longer act fully independently at (e.g., Balogh et al., 2016; Kawinwanichakij et al., 2017; Pintos-Castro et al., 2019). We therefore need to study higher redshift protoclusters to understand how galaxies evolve into their or counterparts.
@@
However, even if it is possible to find such structures, it is difficult to identify which protoclusters are likely to evolve into typical lower redshift clusters. The challenges come from both the sensitivity of cosmological simulations such as Chiang et al. (2013, 2015); Remus et al. (2023); Lim et al. (2024) to different stellar properties as well as the failures to reproduce the high masses and SFR observed in protocluster galaxies. It is thus imperative to ensure that overdensity characterization is as accurate as possible and the properties of their constituent galaxy populations are well constrained, which requires high-quality data and methods to obtain accurate measures of redshift, overdensity strengths, and galaxy properties.
@@
Because protoclusters exist primarily at higher redshifts (), even detecting structure, let alone characterizing a system’s mass, dynamical state, or member population, is highly challenging. In contrast, there have been many successes in finding and characterizing clusters at lower redshifts () from a number of different methodologies. Some notable examples among the highest yielding surveys include the thousands or more candidate clusters found at optical and near-infrared wavelengths (e.g., Gonzalez et al., 2019; Wen & Han, 2022), at X-ray wavelengths (e.g., Ebeling et al., 2001; Piffaretti et al., 2011), and at radio and sub-mm wavelengths by tracing the thermal Sunyaev-Zel’dovich (SZ; Sunyaev & Zeldovich, 1972) effect (e.g., Planck Collaboration et al., 2016; Hilton et al., 2021; Polletta et al., 2021, 2022). Lensing surveys have also seen some success, albeit finding much fewer cluster candidates relative to other methodologies (e.g., Kubo et al., 2009; Ford et al., 2014).
@@
The limiting magnitude required to pick up any meaningful number of galaxies at is out of reach for many optical and near-infrared surveys. This difficulty is compounded for X-ray surveys, as X-ray surface brightness falls off as . Additionally, X-ray and radio/sub-mm surveys, surveys which rely on the identification of signatures from the ICM, are also less effective due to the decreased time such processes are able to act on the member galaxies in order to build up a hot ICM. As a consequence, the ICM has only been detected in very few overdensities relative to the number of searches (e.g., Wang et al., 2016; Tozzi et al., 2022; Di Mascolo et al., 2023). Lensing surveys are susceptible to projection effects complicating the accuracy of mass determination. Red sequence identification is also mostly ineffective due to protocluster galaxies being predominantly blue (Overzier, 2016). Of these shortcomings, the most feasible to overcome is in the optical and near-infrared, where the extreme magnitudes () needed to probe the member population of high-redshift protoclusters is in reach for the deepest ground- and space-based galaxy surveys.
@@
Although the number of known protoclusters at has increased over the past decade due to a number of systematic searches of deep observations (e.g., Diener et al., 2013; Chiang et al., 2014; Wylezalek et al., 2013, 2014; Franck & McGaugh, 2016; Lee et al., 2016; Toshikawa et al., 2016, 2018; Newman et al., 2020; Lee et al., 2024; Ramakrishnan et al., 2024), discovery and characterization has primarily been limited to individual spectacular structures, with some notable examples being the SSA22 protocluster at (Chapman et al., 2001), the Spiderweb protocluster at (Pentericci et al., 1997, 2000; Miley et al., 2006; Dannerbauer et al., 2014), the SPT2349-56 protocluster at (Miller et al., 2018; Hill et al., 2022), and the Hyperion proto-supercluster at (Cucciati et al. 2018, Forrest et al. in prep). The sensitivity offered by the James Webb Space Telescope (JWST) has also enabled protocluster searches beyond , with protocluster candidates being found at redshifts up to (e.g., Laporte et al., 2022; Helton et al., 2024; Li et al., 2024). The focus on a relatively small number of massive protoclusters, however, means that the overall data set is inhomogeneous and thus ill-suited for any broad population analyses, which limits our understanding of the environments and how they affect galaxy evolution at such high redshifts.
@@
In this paper, we perform a search of protoclusters over 2.4 deg2 in three famous extragalactic fields (COSMOS, ECDFS, and CFHTLS-D1). Our search is carried out using a version of Voronoi tessellation Monte-Carlo (VMC, Lemaux et al. 2017; Tomczak et al. 2017; Lemaux et al. 2022) mapping with limited assumptions on the underlying galaxy populations, using as seeds exquisite panchromatic imaging data as well as tens of thousands of spectroscopic redshifts from a variety of archival data sources both public and proprietary. This method has been very successful for systematically finding and characterizing hundreds of new overdensity candidates over 1.4 deg2 between with the Observations of Redshift Evolution in Large-Scale Environments (ORELSE) survey (Hung et al., 2020), including finding clusters of sufficient masses to offer cosmological constraints with only optical and near-infrared data (Hung et al., 2021). With considerable levels of spectroscopy, we are able to effectively search for galaxy-traced high redshift protoclusters over a large range of masses and dynamical states.
@@
Although VMC mapping has also been used to find several massive higher redshift structures in the VIMOS Ultra Deep Survey (e.g., Lemaux et al., 2014a, 2018; Cucciati et al., 2014, 2018; Shen et al., 2021; Forrest et al., 2023; Shah et al., 2024a; Staab et al., 2024), no systematic search has yet been performed within these fields with modern data. Such is the goal of this work. Armed with a wealth of photometric and spectroscopic data, we aim to search for a large ensemble of galaxy-traced protostructures essentially free from any other bias. In this work, we further build on the methodology of Hung et al. (2020) by incorporating a custom-built lightcone with GAlaxy Evolution and Assembly (GAEA; De Lucia et al., 2024) semi-analytics model to estimate the completeness, purity, and masses of the protostructure candidates we find.
@@
This paper is organized as follows. The extensive imaging and spectroscopic data set used is described in §2. The methodology regarding how we find protocluster candidates is described in §3. We describe in §4 how we use simulated data drawn from a custom-built lightcone to quantify the effectiveness of our method at finding and characterizing protocluster candidates. We go over the full catalog of candidates in §5 as well as draw comparisons with the previously known structures in the literature in the fields we studied. Finally, a summary of our findings is given in §6 along with some possible future avenues on expanding upon this work.
@@
Unless otherwise specified, we adopt a flat CDM cosmology throughout this paper, with = 70 km s-1 Mpc-1, = 0.27, and = 0.73, and reported distances are generally given in proper units (e.g., proper kpc or Mpc). In some cases, where explicitly noted, distances are reported in comoving units but still follow the convention of comoving kpc or Mpc.
@@ 2 Data
The VIMOS Ultra Deep Survey (VUDS; Le Fèvre et al., 2015; Tasca et al., 2017) was a spectroscopic redshift survey of roughly 10,000 galaxies at with the goal of studying galaxy assembly. The Observations of Redshift Evolution in Large-Scale Environments (ORELSE; Lubin et al., 2009) survey was a multi-wavelength photometric and spectroscopic campaign designed to map out large-scale structures over . These two surveys were combined into the Charting Cluster Construction with VUDS and ORELSE (C3VO111https://www.orelsesurvey.com/c3vo.html; Shen et al., 2021; Lemaux et al., 2022) survey, which seeks to map out the growth of structure at . The higher redshift end () is an ongoing campaign to obtain additional visible and near-infrared (NIR) wavelength photometry and spectroscopy of three well-studied extragalactic fields: the Cosmic Evolution Survey (COSMOS; Scoville et al., 2007) field, the Extended Chandra Deep Field South (ECDFS; Lehmer et al., 2005), and the first field of the Canada-France-Hawai’i Telescope Legacy Survey (CFHTLS-D1222https://www.cfht.hawaii.edu/Science/CFHTLS/).
@@ 2.1.1 Imaging Data and Photometry
The photometry in the fields at wavelengths relevant to this work generally span 10 or more bands ranging from the ultraviolet (UV) to the NIR. Many bands are extremely deep, reaching magnitudes of for completeness limits of 5.
@@
In the COSMOS field, we utilize the “COSMOS2015” catalog compiled by Laigle et al. (2016) as our basis catalog. The COSMOS2015 catalog consists of near-ultraviolet (NUV) imaging from the GALaxy Evolution eXplorer (GALEX; Martin et al., 2005), ground-based UV/optical imaging from MegaCam on the Canada-France-Hawai’i Telescope (CFHT; Boulade et al., 2003) and Subaru/Suprime-Cam (Miyazaki et al., 2002), ground-based imaging from Subaru/Hyper-Suprime-Cam (HSC; Miyazaki et al., 2012), the UltraVISTA survey (McCracken et al., 2012), and CFHT/WIRCam (Puget et al., 2004; McCracken et al., 2010), as well as 3.6 to 8.0 m imaging from Spitzer with the InfraRed Array Camera (IRAC; Fazio et al., 2004). The catalog contains photometric redshifts for over half a million objects over an area of 2 deg2. We used v2.0 of the Capak et al. (2007) catalog as an additional source for photometric redshifts that were absent in the COSMOS2015 catalog. For the objects shared between the two catalogs, no systematic offset has been observed between the derived parameters for the two different sets of photometry (Lemaux et al., 2022).
@@
Note that while a newer version of the COSMOS photometric catalog exists, namely the COSMOS2020 catalog (Weaver et al., 2022), this catalog did not exist at the inception of this study, and, thus, all mock observations and spectroscopic selection functions used to train and characterize the search algorithm in this study were based on the COSMOS2015 catalog. However, all data in this study are cut at an IRAC channel 1 (hereafter IRAC1) magnitude of IRAC124.8, a limit to which both the COSMOS2015 and COSMOS2020 catalogs are complete. Thus, while there are likely higher order improvements that would come as a result of instead adopting the COSMOS2020 catalog as our base photometric catalog for COSMOS, given that the COSMOS2015 catalog is complete to the depth adopted for this study and the methodological approach used to generate the COSMOS2015 catalog is similar to that of the Classic/Le Phare COSMOS2020 catalog, we chose to keep the COSMOS2015 catalog as our basis for studies in the COSMOS field.
@@
The photometry in the ECDFS field is sourced from optical Subaru/Suprime-Cam imaging over 18 medium bands (Cardamone et al., 2010), taken as a part of the Multiwavelength Survey by Yale-Chile (MUSYC; Gawiser et al., 2006). The field also includes ancillary data by way of imaging from the Garching-Bonn Deep Survey (GaBoDS; Hildebrandt et al., 2006) and imaging from earlier MUSYC observations. Imaging in each of the four Spitzer bands from 3.6 to 8.0 m is also included from the Spitzer IRAC/MUSYC Public Legacy in the ECDF-S (SIMPLE; Damen et al., 2011) survey. These data overall include over 40,000 photometric redshifts over a 0.25 deg2 area. We note that while this field also has observations taken with the Hubble Space Telescope (HST) for the Cosmic Assembly Near-infrared Deep Extragalactic Legacy Survey (CANDELS; Grogin et al., 2011; Koekemoer et al., 2011), the area observed only intersects a small section of the spectroscopic coverage.
@@
The CFHTLS-D1 field’s UV/optical/NIR imaging data is described extensively in Lemaux et al. (2014b, a). As part of CFHTLS, the data include ground-based imaging with CFHT/MegaCam. In the near-infrared, imaging was taken with CFHT/WIRCam for the WIRCam Deep Survey (WIDS; Bielby et al., 2012) and at 3.6 and 4.5 m with Spitzer/IRAC for the Spitzer Extragalactic Representative Volume Survey (SERVS; Mauduit et al., 2012). The field includes over 150,000 photometric redshifts over a 1 deg2 area.

### v2

@@ Abstract
The Charting Cluster Construction with VUDS and ORELSE (C3VO) survey is an ongoing imaging and spectroscopic campaign aiming to map out the growth of structure up to and was born from the combination of the Visible Multi-Object Spectrograph Ultra Deep Survey and the Observations of Redshift Evolution in Large-Scale Environments (ORELSE) survey. As we previously accomplished with the ORELSE survey, we apply our technique known as Voronoi tessellation Monte Carlo (VMC) mapping to search for serendipitous galaxy overdensities at in the three C3VO fields. We also apply the same technique to mock observations of simulated galaxies with properties derived from the GAlaxy Evolution and Assembly semianalytic model in order to judge the effectiveness of our search algorithm as a function of redshift, total mass, and fraction of spectroscopic redshifts. We find completeness and purity values of the order of 30-50% for and , with a strong dependence on mass and redshift, with values as high as 80% and 70%, respectively, in the best-case scenario for . In the C3VO fields, we were able to recover many of the previously known structures in the literature as well as find hundreds of new overdensity candidates, once again demonstrating the powerful capabilities of VMC mapping when applied to wide-field optical and infrared galaxy evolution surveys at ever higher redshifts.
@@ 1 Introduction
Galaxy groups and clusters are among the largest gravitationally bound structures in the Universe. These large-scale structures (LSSs) at different redshifts provide a means of studying cosmology as well as the effects of environment on galaxy evolution at various stages in time. Protoclusters are the progenitors of galaxy clusters, which, at some point, will gravitationally collapse and become virialized. Clusters may be distinguished from field galaxies through a number of methods, such as the identification of a well-defined red sequence or a hot intracluster medium (ICM). Red sequence techniques to find clusters were first employed by the Red Sequence Cluster Survey (RCS; Gladders & Yee, 2000, 2005). The RCS made use of two filters, one in the optical and one in the near-infrared (NIR), to identify galaxy overdensities by features such as Balmer absorption or the 4000Å break, which would be present in the filter wavelengths out to redshifts of . With the addition of mid-infrared filters, red sequence searches can be extended up (e.g., Andreon et al., 2009; Muzzin et al., 2009; Wilson et al., 2009; Galametz et al., 2012; Cooke et al., 2015; Noirot et al., 2016; Golden-Marx et al., 2019).
@@
Various cluster surveys at intermediate redshift, such as the Observations of Redshift Evolution in Large-Scale Environments (ORELSE; Lubin et al., 2009) survey, Gemini Cluster Astrophysics Spectroscopic Survey (GCLASS; Muzzin et al., 2012), and the Gemini Observations of Galaxies in Rich Early ENvironments (GOGREEN; van der Burg et al., 2020) survey, as well as earlier cluster studies (e.g., Aragon-Salamanca et al., 1993; Kodama et al., 1998; Lubin et al., 1998; Stanford et al., 1998), have demonstrated how the scaling relations between color, density, morphology, and the star formation rate (SFR) change up to relative to the local Universe. Studies from these surveys have shown that quenching is likely driven by internal mass-based mechanisms as well as processes that are at play in overdense environments. There are claims that the effects of these processes are largely separable for galaxy samples at (e.g., Peng et al., 2010; Kovač et al., 2014; Guglielmo et al., 2015), but some studies argue otherwise (e.g., De Lucia et al., 2012). The scaling relations weaken or reverse as redshift increases, and the two modes of quenching may also no longer act fully independently at (e.g., Balogh et al., 2016; Kawinwanichakij et al., 2017; Pintos-Castro et al., 2019). We therefore need to study higher-redshift protoclusters to understand how galaxies evolve into their or counterparts.
@@
However, even if it is possible to find such structures, it is difficult to identify which protoclusters are likely to evolve into typical lower-redshift clusters. The challenges come from both the sensitivity of cosmological simulations such as Chiang et al. (2013, 2015), Remus et al. (2023), and Lim et al. (2024) to different stellar properties as well as the failures to reproduce the high masses and SFR observed in protocluster galaxies. It is thus imperative to ensure that overdensity characterization is as accurate as possible and the properties of their constituent galaxy populations are well constrained, which requires high-quality data and methods to obtain accurate measures of redshift, overdensity strengths, and galaxy properties.
@@
Because protoclusters exist primarily at higher redshifts (), even detecting structure with ground-based telescopes, let alone characterizing a system’s mass, dynamical state, or member population, is highly challenging. In contrast, there have been many successes in finding and characterizing clusters at lower redshifts () from a number of different methodologies. Some notable examples among the highest-yielding surveys include the thousands or more candidate clusters found at optical and NIR wavelengths (e.g., Gonzalez et al., 2019; Wen & Han, 2022), at X-ray wavelengths (e.g., Ebeling et al., 2001; Piffaretti et al., 2011), and at radio and submillimeter wavelengths by tracing the thermal Sunyaev-Zel’dovich (SZ; Sunyaev & Zeldovich, 1972) effect (e.g., Planck Collaboration et al., 2016; Hilton et al., 2021; Polletta et al., 2021, 2022). Lensing surveys have also seen some success, albeit finding far fewer cluster candidates relative to other methodologies (e.g., Kubo et al., 2009; Ford et al., 2014).
@@
The limiting magnitude required to pick up any meaningful number of galaxies at is out of reach for many optical and NIR surveys (e.g., a limiting magnitude of in the band probes 0.3 over ). This difficulty is compounded for X-ray surveys, as X-ray surface brightness falls off as . Additionally, X-ray and radio/submillimeter surveys (surveys which rely on the identification of signatures from the ICM) are also less effective due to the decreased time that such processes are able to act on the member galaxies in order to build up a hot ICM. As a consequence, the ICM has only been detected in very few overdensities relative to the number of searches (e.g., Wang et al., 2016; Tozzi et al., 2022; Di Mascolo et al., 2023). Lensing surveys are susceptible to projection effects complicating the accuracy of mass determination. Red sequence identification is also mostly ineffective due to protocluster galaxies being predominantly blue (Overzier, 2016). Of these shortcomings, the most feasible to overcome is in the optical and NIR, where the extreme magnitudes () needed to spectroscopically probe the member population of high-redshift protoclusters over a large area (i.e., tens of arcminutes on the sky) are within reach for the deepest ground- and space-based galaxy surveys.
@@
Although the number of known protoclusters at has increased over the past decade due to a number of systematic searches of deep observations (e.g., Diener et al., 2013; Chiang et al., 2014; Wylezalek et al., 2013, 2014; Franck & McGaugh, 2016; Lee et al., 2016; Toshikawa et al., 2016, 2018; Newman et al., 2020; Lee et al., 2024; Ramakrishnan et al., 2024), discovery and characterization have primarily been limited to individual spectacular structures, with some notable examples being the SSA22 protocluster at (Chapman et al., 2001), the Spiderweb protocluster at (Pentericci et al., 1997, 2000; Miley et al., 2006; Dannerbauer et al., 2014), the SPT2349-56 protocluster at (Miller et al., 2018; Hill et al., 2022), and the Hyperion proto-supercluster at (Cucciati et al. 2018, Forrest et al. 2025, in preparation). The sensitivity offered by the James Webb Space Telescope (JWST) has also enabled protocluster searches beyond , with protocluster candidates being found at redshifts up to (e.g., Laporte et al., 2022; Helton et al., 2024; Li et al., 2024). The focus on a relatively small number of massive protoclusters, however, means that the overall data set is inhomogeneous and thus ill-suited for any broad population analyses, which limits our understanding of the environments and how they affect galaxy evolution at such high redshifts.
@@
In this paper, we perform a search of protoclusters over 2.4 deg2 in three well-studied extragalactic fields (COSMOS, ECDFS, and CFHTLS-D1). Our search is carried out using a version of Voronoi tessellation Monte Carlo (VMC; Lemaux et al., 2017; Tomczak et al., 2017; Lemaux et al., 2022) mapping with limited assumptions on the underlying galaxy populations, using as seeds exquisite panchromatic imaging data as well as tens of thousands of spectroscopic redshifts from a variety of archival data sources both public and proprietary. This method has been very successful for systematically finding and characterizing hundreds of new overdensity candidates over 1.4 deg2 between with the ORELSE survey (Hung et al., 2020), including finding clusters of sufficient masses to offer cosmological constraints with only optical and NIR data (Hung et al., 2021). Our past successes have shown that, with considerable levels of spectroscopy (see Hung et al. 2020 for a through breakdown per field), we are able to effectively search for galaxy-traced high-redshift protoclusters over a large range of masses and dynamical states.
@@
Although VMC mapping has also been used to find several massive higher-redshift structures in the Visible Multi-Object Spectrograph (VIMOS) Ultra Deep Survey (e.g., Lemaux et al., 2014a, 2018; Cucciati et al., 2014, 2018; Shen et al., 2021; Forrest et al., 2023; Shah et al., 2024a; Staab et al., 2024), no systematic search has yet been performed within these fields. Such is the goal of this work. Armed with tens of thousands of photometric and spectroscopic redshifts, we aim to search for a large ensemble of galaxy-traced protostructures essentially free from any other bias. In this work, we further build on the methodology of Hung et al. (2020) by incorporating a custom-built lightcone with a GAlaxy Evolution and Assembly (GAEA; De Lucia et al., 2024) semianalytics model to estimate the completeness, purity, and masses of the protostructure candidates we find.
@@
This paper is organized as follows. The extensive imaging and spectroscopic data set used is described in Section 2. The methodology regarding how we find protocluster candidates is described in Section 3. We describe in Section 4 how we use simulated data drawn from a custom-built lightcone to quantify the effectiveness of our method at finding and characterizing protocluster candidates. We go over the full catalog of candidates in Section 5 as well as draw comparisons with the previously known structures in the literature in the fields we studied. Finally, a summary of our findings is given in Section 6 along with some possible future avenues on expanding upon this work.
@@
Unless otherwise specified, we adopt a flat cold dark matter (CDM) cosmology throughout this paper, with = 70 km s-1 Mpc-1, = 0.27, and = 0.73, and reported distances are generally given in proper units (e.g., proper kiloparsecs or megaparsecs). In some cases, where explicitly noted, distances are reported in comoving units but still follow the convention of comoving kiloparsecs or megaparsecs.
@@ 2 Data
The VIMOS Ultra Deep Survey (VUDS; Le Fèvre et al., 2015; Tasca et al., 2017) was a spectroscopic redshift survey of roughly 10,000 galaxies at with the goal of studying galaxy assembly. The ORELSE (Lubin et al., 2009) survey was a multiwavelength photometric and spectroscopic campaign designed to map out LSSs over . These two surveys were combined into the Charting Cluster Construction with VUDS and ORELSE (C3VO111https://www.orelsesurvey.com/c3vo.html; Shen et al., 2021; Lemaux et al., 2022) survey, which seeks to map out the growth of structure at . The higher-redshift end () is an ongoing campaign to obtain additional visible and NIR wavelength photometry and spectroscopy of three well-studied extragalactic fields: the Cosmic Evolution Survey (COSMOS; Scoville et al., 2007) field, the Extended Chandra Deep Field South (ECDFS; Lehmer et al., 2005), and the first field of the Canada–France–Hawai’i Telescope Legacy Survey (CFHTLS-D1222https://www.cfht.hawaii.edu/Science/CFHTLS/).
@@ 2.1.1 Imaging Data and Photometry
The photometry in the fields at wavelengths relevant to this work generally span 10 or more bands ranging from the ultraviolet (UV) to the NIR. Many bands are extremely deep, reaching magnitudes of for completeness limits of 5. We provide a brief summary of the archival data used in this work below, but we refer the reader to Laigle et al. (2016), Cardamone et al. (2010), and Lemaux et al. (2014a) and the references therein for more details on the COSMOS, ECDFS, and CFHTLS-D1 fields, respectively.
@@
In the COSMOS field, we utilize the “COSMOS2015” catalog compiled by Laigle et al. (2016) as our basis catalog. The COSMOS2015 catalog consists of near-UV imaging from the GALaxy Evolution eXplorer (GALEX; Martin et al., 2005), ground-based UV/optical imaging from MegaCam on the Canada–France–Hawai’i Telescope (CFHT; Boulade et al., 2003) and Subaru/Suprime-Cam (Miyazaki et al., 2002), ground-based imaging from Subaru/Hyper-Suprime-Cam (Miyazaki et al., 2012), the UltraVISTA survey (McCracken et al., 2012), and CFHT/WIRCam (Puget et al., 2004; McCracken et al., 2010), as well as 3.6–8.0 m imaging from Spitzer with the InfraRed Array Camera (IRAC; Fazio et al., 2004). The catalog contains photometric redshifts for over half a million objects over an area of 2 deg2. We used v2.0 of the Capak et al. (2007) catalog as an additional source for photometric redshifts that were absent in the COSMOS2015 catalog. For the objects shared between the two catalogs, no systematic offset has been observed between the derived parameters for the two different sets of photometry (Lemaux et al., 2022).
@@
Note that while a newer version of the COSMOS photometric catalog exists, namely the COSMOS2020 catalog (Weaver et al., 2022), this catalog did not exist at the inception of this study, and, thus, all mock observations and spectroscopic selection functions used to train and characterize the search algorithm in this study were based on the COSMOS2015 catalog. However, all data in this study are cut at an IRAC channel 1 (hereafter IRAC1) magnitude of IRAC124.8, a limit to which both the COSMOS2015 and COSMOS2020 catalogs are complete. While there are likely higher-order improvements that would come as a result of instead adopting the COSMOS2020 catalog as our base photometric catalog for COSMOS, we do not expect its usage to alter our findings. Moreover, given that the COSMOS2015 catalog is complete to the depth adopted for this study and the methodological approach used to generate the COSMOS2015 catalog is similar to that of the Classic/LE PHARE COSMOS2020 catalog, we chose to keep the COSMOS2015 catalog as our basis for studies in the COSMOS field.
@@
The photometry in the ECDFS field is sourced from optical Subaru/Suprime-Cam imaging over 18 medium bands (Cardamone et al., 2010), taken as a part of the Multiwavelength Survey by Yale-Chile (MUSYC; Gawiser et al., 2006). The field also includes ancillary data by way of imaging from the Garching-Bonn Deep Survey (GaBoDS; Hildebrandt et al., 2006) and imaging from earlier MUSYC observations. Imaging in each of the four Spitzer bands from 3.6–8.0 m is also included from the Spitzer IRAC/MUSYC Public Legacy in the ECDF-S (SIMPLE; Damen et al., 2011) survey. These data overall include over 40,000 photometric redshifts over a 0.25 deg2 area. We note that while this field also has observations taken with the Hubble Space Telescope (HST) for the Cosmic Assembly Near-infrared Deep Extragalactic Legacy Survey (CANDELS; Grogin et al., 2011; Koekemoer et al., 2011), the area observed only intersects a small section of the spectroscopic coverage.
@@
The CFHTLS-D1 field’s UV/optical/NIR imaging data is described extensively in Lemaux et al. (2014b, a). As part of CFHTLS, the data include ground-based imaging with CFHT/MegaCam. In NIR, imaging was taken with CFHT/WIRCam for the WIRCam Deep Survey (WIDS; Bielby et al., 2012) and at 3.6 and 4.5 m with Spitzer/IRAC for the Spitzer Extragalactic Representative Volume Survey (Mauduit et al., 2012). The field includes over 150,000 photometric redshifts over a 1 deg2 area.